Thomas Brüstle: On tame cluster algebras

Cluster algebras have been introduced by Fomin and Zelevinsky in 2001. Since then, the cluster algebras of finite type have been classified, and various connections to tilted algebras, quadratic forms and their corresponding finite type criteria have been found. We adress in this talk the question how tame cluster algebras can be defined and characterized.